### v2

# Time Resolved Absorption of Six Chemical Species With MAROON-X Points to a Strong Drag in the Ultra Hot Jupiter TOI-1518 b

Abstract
Context. Wind dynamics play a pivotal role in governing transport processes within planetary atmospheres, influencing atmospheric chemistry, cloud formation, and the overall energy budget. Understanding the strength and patterns of winds is crucial for comprehensive insights into the physics of ultra-hot Jupiter atmospheres. Current research has proposed two contrasting mechanisms that limit wind speeds in these atmospheres, each predicting a different scaling of wind speed with planet temperature. However, the sparse nature of existing observations hinders the determination of population trends and the validation of these proposed mechanisms.
Aims. This study focuses on unraveling the wind dynamics and the chemical composition in the atmosphere of the ultra-hot Jupiter TOI-1518 b
Methods. Two transit observations using the high-resolution ( 000), optical (spectral coverage between 490 and 920 nm) spectrograph MAROON-X were obtained and analyzed to explore the chemical composition and wind dynamics using the cross-correlation techniques, global circulating models, and atmospheric retrieval.
Results. We report the detection of 14 species in the atmosphere of TOI-1518 b through cross-correlation analysis. Additionally, we measure the time-varying cross-correlation trails for 6 different species, compare them with predictions from General Circulation Models (GCM) and conclude that a strong drag is present in TOI-1518b’s atmosphere (). The ionized species require stronger drags than neutral species, likely due to the increased magnetic effects in the upper atmosphere. Furthermore, we detect vanadium oxide (VO) using the most up-to-date line list. This result is promising in detecting VO in other systems where inaccuracies in previous line lists have hindered detection. We use a retrieval analysis to further characterize the abundances of the different species detected. Chromium, magnesium, vanadium, and titanium are found with lower than solar abundances, possibly due to the ionization of these elements or to their incorporation into TiO and VO.
1 Introduction
The golden age of exoplanet characterisation began in the last two decades. One of the most exciting topics is the exploration of their atmospheric diversity in terms of composition and dynamics
(Madhusudhan, 2019; Wordsworth & Kreidberg, 2022). Exoplanet atmosheres can be studied by observing their spectra either in emission (Chauvin et al., 2005; Swain et al., 2008), in transmission (Charbonneau et al., 2002) and soon in reflected light (Martins et al., 2013) from ground-based and space observatories. Recently, the sensitivity of
the JWST has begun to revolutionise this field, enabling extremely advanced studies of the fine structure and dynamics of the atmospheres of giant planets (Tsai et al., 2023; Coulombe et al., 2023). Thanks to their extended atmospheres, ultra-hot Jupiters (UHJs) are ideal targets for atmospheric characterization in transmission. These planets are very close to their stars and are tidally locked. This implies a significant day/night temperature gradient, which creates strong atmospheric circulation (Showman et al., 2020). Differences in temperature of several hundred degrees have been measured between the daysides and nightsides (Parmentier & Crossfield, 2018).
Due to their extreme temperature, volatile and refractory elements are accessible and detectable in such atmospheres. Indeed refractory species (with high condensation temperature) are expected to be gaseous in UHJs (Lothringer et al., 2018) when, in colder planets they are inaccessible because they condensed out of the gas phase. The measure of the refractory to volatile elemental ratio of these planets (e.g., O/Fe. C/Fe), recently emerged as a new powerful way to trace planet formation (Lothringer et al., 2021; Chachan et al., 2023; Pelletier et al., 2024; Smith et al., 2024).
When using low to moderate spectral resolution (), like that provided by a space telescope such as JWST, the spectra captured contain a mixture of information from various parts of the atmosphere. Each part of the atmosphere has different properties, such as temperature or chemical composition (Espinoza et al., 2024). This information mixture could lead to misleading or biased inferences about the different properties derived from the data (Feng et al., 2015; Line & Parmentier, 2016).
For the first time in 2010, high-resolution () spectroscopy was used to characterize the atmosphere of a transiting ultra-hot Jupiter by resolving individual molecular lines using CRIRES at VLT (Snellen et al., 2010). During the transit, the Doppler shifts caused by the planet’s rotation and the atmospheric winds allow lines formed in different parts of the planetary atmosphere to be spectroscopically separated (see Nortmann et al. 2024 for an extreme example). Recent ESPRESSO observations at VLT showed that the iron absorption lines of WASP-76b and WASP-121b, two canonical Ultra-Hot Jupiters (UHJs), are progressively blueshifted during the transit (Ehrenreich et al., 2020; Borsa et al., 2021).
While two competing scenarios have been suggested to explain this behavior, the precise physical mechanism remains elusive.
The signal could result from a hot, puffy evening terminator and a cool, compact morning terminator, whereby the blueshifting winds and rotation of the hot evening terminator dominate the absorption signal due to its larger-scale height (Wardenier et al., 2021). Alternatively, it has been shown that 3D models with opaque clouds can also reproduce the observed signal (Savel et al., 2022), whereby the cloud deck “blocks” the absorption features on the morning terminator. Other studies highlights the richness of WASP-76b with the detection of multiple species with different shift (Kesseli et al., 2022).
To get more insights into this wind speed problem, we observed the UHJ TOI-1518 b. This planet is orbiting a fast-rotating F0-type star of 7300K. Table 1 summarizes the physical properties of TOI-1518 and TOI-1518 b. This planet (T = 2498 K) is ideal for a direct comparison with the well-studied UHJs WASP-76 b (T = 2228 K) and WASP-121 b (T = 2720 K) with iron previously detected by Cabot et al. (2021).
In this paper, we present the observations of two transits of TOI-1518 b observed with MAROON-X at the Gemini-North Observatory. After presenting the observations and the data reduction in Sect. 2, we present the chemical information we obtained thanks to Cross-Correlation techniques in Sect. 3. Then, we compare the iron trail detected with global circulating models (GCMs) to explore the wind dynamics of the planet in Sect. 4. We finally present a retrieval analysis in Sect. 5, hinting at the different abundances of the species detected in the atmosphere of TOI-1518 b.
| Stellar parameters | Value | Ref |
| Stellar radius | [1] | |
| Effective temperature | [1] | |
| Metallicity [Fe/H] | [1] | |
| Rotational velocity | [1] | |
| Spectral type | F0 | [1] |
| Planetary parameters | Value | Ref |
| Planet mass | [1] | |
| Planet radius | [1] | |
| Equilibrium temperature | [1] | |
| System parameters | Value | Ref |
| Orbital period | [1] | |
| Mid-transit Time | [1] | |
| Orbital Inclination | [1] | |
| Semi-major axis | [1] | |
| Systemic Velocity () | [1] | |
| Projected orbital velocity () | ||
| Impact Parameter (b) | [1] |
2 Observations and Data reduction
Two transits of the UHJ TOI-1518 b were observed with MAROON-X, a high-resolution () optical (spectral coverage between 490 and 920 nm) spectrograph at the 8.1-m Gemini-North observatory in Hawaii. Recent observations showed the capacity of MAROON-X to characterize UHJs by detecting ions and volatile and refractory elements. It has also allowed the description of the trail of the atmospheric signal (Pelletier et al., 2023; Prinoth et al., 2023) and even to study some strong lines such as Ca+ triplet (Prinoth et al., 2024). The observations were taken on 2022-08-13 and 2023-10-19 (program ID GN-2022B-Q-128 and GN-2023B-Q-127, PI: Parmentier). A summary of the observations is given in Table 2.
| Night | 2022-08-13 | 2023-10-19 |
| Phase | 0.96-0.04 | 0.96-0.04 |
| Nobs | 40 (25 + 15) | 41 (25 + 16) |
| Exp. time | 260s (b), 220s (r) | 220s (both arms) |
| Airmass | 1.45 - 1.9 | 1.47 - 1.6 |
| S/N | 110 - 180 | 155 - 210 |
| (Avg = 150) | (Avg = 180) |
Notes. Nobs is the total number of observed spectra with in-transit (25) and out-of-transit (15) observations
2.1 Data Quality
MAROON-X is divided into two detectors, one ”blue” covering wavelengths ranging from 490 to 678 nm. The second one, ”red,” covers wavelengths ranging from 640 to 920 nm. To ensure complete coverage of transit events, the observation protocol includes baseline measurements both pre- and post-transit. The exposure time for the red detector was slightly lower than for the blue arm (220s vs 260s) during the first observation. For the second observation, the exposure time for both detectors was set up to 220 sec. Each order of the red detector comprises 4036 pixels, while the blue one has only 3954 pixels. There are 28 spectral orders for the red detector and 33 for the blue one. Fig 1 presents the signal-to-noise ratio (S/N) and the airmass as a function of the observation frames for both transits. For both nights, the S/N was at least above 110, with a slightly better S/N for the second night (155 min and 210 max). During the second night, the blue arm seemed less performant than the red arm, but this is because the blue arm’s exposure time was higher during the first transit. Conditions of the second transit (average humidity = 7%) were so good that both arms still had better S/N than during the first transit (average humidity =24%).
The MAROON-X data were reduced using the standard pipeline (Seifahrt et al., 2020) in one-dimensional wavelength-calibrated spectra, order by order for each time series exposure. The outputs are given as with the number of spectral orders. In total, 40 frames were observed during both transits. The redder order of the blue detector (between 668 and 678nm) has been removed because of a too-low S/N (¡ 35).
One main limitation with high-resolution transmission spectroscopy from ground-based observations is our atmosphere and stellar signals. Planetary signals are much fainter but change over time because of the rapid Doppler acceleration, inducing shifts of many tens of km.s-1 to the planet spectrum over the transit duration. The telluric lines stay constant, and the positions of the stellar lines vary in order of hundreds of m.s-1. This is why we can disentangle the planetary signal from the others. We applied different reduction steps on the data following the sequence described by Pelletier et al. (2023) and summarized below:
All observed spectra are aligned in the stellar rest frame to remove the Earth’s barycentric motion and TOI-1518’s reflex motion. This is necessary to subtract the stellar signal from the data.
Each spectrum is set to the same continuum level to remove blaze and throughput variations.
The in-transit data are divided by a master stellar spectrum made with the out-of-transit data.
A principal component analysis (PCA) approach removes the telluric signal and residuals from the stellar correction. We selected three number of principal components to remove from each species. Higher number of components would start affecting the planetary atmospheric signal. The Ca+ triplet is so strong that no PCA is needed to detect it, as shown in Fig. 2. Both panels show two orders of MAROON-X red detectors where the lines of the Ca+ triplet are located. The top panel shows the data after pipeline extraction with the strong Ca+ stellar lines. The lower panel shows the residuals of the data reduction that preceded the PCA. The strong stellar Ca+ lines are removed using the out-of-transit data, leaving the planetary signal with the Doppler shadow effect (see sub-Section 3.3). After this step, the cleaned data are ready for Cross-Correlation to study the atmospheric composition of TOI-1518 b.
3 Cross-Correlation analysis
3.1 Template spectra for cross-correlation
The Cross-Correlation method is needed to detect the faint planetary lines in the residuals obtained after PCA. We cannot detect most of them directly except for a few individual lines (such as Ca+). Fortunately, atoms and molecules have so many spectral lines that combining the signal with CCF boosts the S/N, allowing a precise detection of the planetary signal. A template is needed to combine these lines. Synthetic spectra of TOI-1518 b have been generated with PetitRadtrans (Mollière et al., 2019) and FastChem (Stock et al., 2022), which respect equilibrium chemistry.
The spectra for individual molecules were produced at a resolution of over the 400 to 1000 nm wavelength range. The input parameters are those of the TOI-1518 system as shown in Table. 1. Collision induced absorption (CIA) cross-sections for the H2-H2 and H2-He pairs are used to generate a continuum opacity.
Subsequently, these spectra have been interpolated onto the MAROON-X wavelength grid and convolved to match the instrumental resolution.
The spectra were also convolved with the planetary rotation rate. These spectra are shown in Fig. 3. The species selected in this study are based on those previously detected in recent MAROON-X publications (Prinoth et al., 2023; Pelletier et al., 2023).
Line list from Kurucz database were used for all atoms and ions (Kurucz, 2017). For TiO, we used TOTO line list (McKemmish et al., 2019). For VO, both the HyVO line list (Bowesman et al., 2024) and the VOmyt line list (McKemmish et al., 2016) outputs were compared in Sect.3.2 and 3.4.
Alkaline metals and ions show individual solid lines, while other metals show line forests. Most of the signals except for the Ca+ triplet are stronger in the blue part of the spectrum, which corresponds to the blue detector of MAROON-X. We decided to analyze each detector individually as if it were two different transits and then to sum every - map or CCF map.
3.2 Species detected in the - maps
The cross-correlation maps are converted into velocity-velocity maps (- diagrams) by shifting them towards the expected rest-frame of the planet ( = 218 6.79 km.s-1, = 0.17 km.s-1), assuming values of projected orbital velocity between 0 and 400 km.s-1 in steps of 1 km. s-1. Fig. 4 shows fourteen clear detections obtained thanks to the CCF. The white cross in each plot is the position of the expected signal from the planet if the planetary atmosphere is considered static and the planet has a circular orbit. The signal is in the rest frame of the planet due to the previous correction of the reduction process from the systemic and barycentric velocity. The expected is calculated as follows:
with , where is the period, is the semi-major axis of TOI-1518 b and the inclination of the system. All these parameters are given in Table 1. To compute our - plots, we selected a range of orbital velocity () from 0 to 400 km.s-1 and a range of rest-frame velocity () from -150 to 150 km.s-1 with steps of 1 km.s-1 for each. We then integrate each point of the CCF maps previously obtained (see Fig. 4) following the slope determined by the orbital velocity at the rest frame position determined by . Fourteen species are detected with an S/N 4.5. The noise level is calculated in a region far from the central peak at 75 km.s-1 where no signal of the planet or Rossiter-McLaughlin (RM) residuals is expected. The parameters of the best Gaussian fits of the maps are presented in Table 3.
| Kp | Amp () | V | FWHM | |
| Ca+ | -84.8 0.5 | 21.1 0.5 | 4.9 0.1 | 36.9 0.2 |
| Fe+ | -17.8 0.1 | 12.3 1.1 | -4.0 0.5 | 11.4 1.2 |
| Ti+ | -26.7 0.1 | 7.5 1.1 | -3.4 0.6 | 9.1 1.4 |
| Ba+ | -26.6 0.1 | 11.4 -0.6 | -5.2 0.3 | 13.7 0.8 |
| Ca | -22.1 0.1 | 8.1 0.9 | -3.7 0.6 | 10.1 1.4 |
| Fe | -32.9 0.2 | 17.9 1.7 | -3.8 0.5 | 11.1 1.2 |
| Ti | -15.2 0.4 | 10.1 0.7 | -4.3 0.3 | 9.7 0.8 |
| Si | -39.9 0.2 | 5.9 0.9 | -1.4 0.8 | 9.7 1.8 |
| Mg | -0.8 0.8 | 8.1 0.8 | -3.1 0.6 | 11.7 1.4 |
| V | -58.1 0.6 | 5.8 0.5 | -3.1 0.5 | 10.2 1.1 |
| Mn | -16.2 0.6 | 4.0 0.2 | -3.5 0.2 | 12.7 0.6 |
| Cr | -19.5 0.1 | 4.0 0.4 | -1.3 0.8 | 12.5 1.9 |
| Na | +13.4 2.0 | 7.9 1.0 | -1.3 1.2 | 18.3 3.3 |
| VO | -43.5 0.2 | 4.9 0.2 | -12.1 0.2 | 10.5 0.5 |
| All | -31.2 0.1 | 21.1 2.6 | -2.9 0.7 | 11.6 1.7 |
The of other species of interested detected in other UHJs are presented in Appendix (see Fig 17). These ones show positive correlation near the expected orbital position and may warrant follow-up observations. For some species, such as iron, calcium, magnesium, or titanium, the signal is blueshifted compared to the expected velocity (Vres = km.s-1). The orbital velocity is also lower than expected (km.s-1).
3.3 Trails of the signal in CCF maps
TOI 1518 b is a misaligned planet, as presented in Cabot et al. (2021). A nodal precession analysis made by Watanabe et al. (2024) underline a change of the impact parameter of the planet in function of time (from b = 0.91497 in 2019 to b= 0.8797 in 2022). However, this change does not affect significantly our result and the geometry of the system in our short timescale so we decided to keep the value of Cabot et al. (2021) (b=0,9036) as reference for our study. The geometry of the system is presented in Fig. 16. Fig. 5 presents the iron cross-correlation map on the stellar rest frame on the left and shifted to the planetary rest frame on the right. The dashed yellow line on the right underlines the expected position of the planet if the atmosphere is static. As the planet is misaligned, the Doppler shadow effect (already observed in Cont et al. 2021) only affects the planetary signal at the beginning of the transit and we have decided to not mask it or remove it. This is a different case from WASP-76 b (Pelletier et al., 2023; Ehrenreich et al., 2020) where it has to be removed as the Doppler shadow crosses the planetary trail.
The iron track shown in Fig. 5 is slightly shifted from the theoretical planetary velocity computed with the orbital parameters of Table 2. Previous high-resolution observations have demonstrated the ability to resolve time variations of this atmospheric track (Ehrenreich et al., 2020; Borsa et al., 2021). We thus investigated this with our data. We binned the CCF similarly to Wardenier et al. (2024) to increase the planetary signal. We divided the 25 in-transit frames observed in both datasets into nine bins. We use scipy.optimize.curve fit to fit a Gaussian to each bin between 10 km/s (in the planetary rest frame). This gave better results than performing the fit across a broader range of velocities. The results are shown in Fig. 6. The velocity center of the atmospheric track changes with time (Fig. 6, Left Panel), becoming more blue-shifted from around +1 km.s-1 to around -8 km.s-1. This is similar to what happens in the case of other UHJs like WASP-76b and WASP-121b (Ehrenreich et al., 2020; Borsa et al., 2021). The signal amplitude varies, with the function of the orbital phase being more important at mid-transit than at the beginning or end of the transit.
We further detect the time-varying trace of the CCF for six different species: Fe, Fe+, Ca, Ca+, Na, Mg (Fig. 7). All species show a qualitatively similar behavior to the Fe trace, with a blueshift over time.
3.4 Discussion
The equilibrium temperature of TOI-1518 b (2492 K, Cabot et al. 2021) is lower than WASP-189 b (2641 K, Anderson et al. 2018) and higher than WASP-76 b (2228 K, Ehrenreich et al. 2020), two recently UHJs observed with MAROON-X (Pelletier et al., 2023; Prinoth et al., 2023). Then, each species observed in both planets is expected to be present in TOI-1518 b. The detection of iron, manganese, chromium, vanadium, magnesium, calcium, and sodium is thus consistent with previous observations. The non-detection of potassium is due to the overlap of the telluric water lines with the Doppler-shifted potassium lines. This is the consequence of an unfortunate systemic velocity and barycentric velocity during these two observations. The detection of titanium in TOI-1518 b, while it was not present on the cooler WASP-76 b but present on the hotter WASP-189 b, might be a sign that there is a trend of titanium abundance with temperature, possibly linked to the formation of TiO or to the nightside condensation of titanium. Whereas TiO would be expected in TOI-1518b, we are not able to find it.
The detection of VO was made with a significance of 4.9 sigma using the newly released HyVO line list this year (Bowesman et al., 2024). VO is notoriously difficult to detect in exoplanet atmospheres; however, Pelletier et al. 2023 demonstrates the feasibility of such detections when employing a more accurate line list. For WASP-76b, the VOmyt line list was useful for the detection of VO. Fig. 8 illustrates that this line list was unsuccessful in retrieving VO signals in the observations of TOI-1518b. This discrepancy can be attributed to a stronger VO signal in WASP-76b, along with the use of three transit observations compared to only two for TOI-1518b. In this study, we demonstrate the superior capability of the HyVO line list, which successfully detected VO (see Fig. 8). The detected signal is blueshifted by as much as -12.11 0.22 km/s (see Table 3), which is significantly higher than the maximum blueshift observed for other species in TOI-1518b, recorded at -5.16 0.57 km/s. Currently, it remains uncertain whether this difference is due to a shift in the line list itself or a physical shift in the atmosphere, potentially stemming from the varying localization of VO compared to metals and ions in the atmosphere of TOI-1518b. The presence of strong optical absorbers such as VO and TiO in the atmosphere of UHJs is crucial for thermal inversion phenomena (Fortney et al., 2008). This finding underscores the significance of utilizing more accurate line lists for the detection of these molecules, especially in cases where the signals may be weaker than those observed in WASP-76b. Further studies employing this new line list could reveal the presence of VO in targets where it was previously undetectable, as highlighted in Borsa et al. (2021).
Detecting Fe/Fe+, Ca/Ca+, Ti/Ti+, and potentially V/VO raises questions. Due to their presence in specific parts of the atmosphere, these species should not co-exist. For example, Fe+ is expected more in the hotter dayside atmosphere, whereas Fe should be more present on the cooler limbs and nightside. However, the tracks of Fe and Fe+ seen in Figure 7 show a similar blueshifting trend, meaning that they are likely probing similar atmospheric regions, albeit probably at different pressures.
Additionally, the iron trail of TOI-1518 b observed in Fig. 5 and Fig. 6 follows a comparable trend than the ones previously observed in WASP-76 b (Ehrenreich et al., 2020), WASP-121 b (Borsa et al., 2021) and WASP-189 b (Prinoth et al., 2023) with a signal becoming more blueshifted with the transit.
One of the main uncertainties of this work is the impact of the PCA step on the planetary signal. Fig. 22 shows the iron map for three components removed on the left, and on the right, it shows the maximum of the map in the function of the number of PCA components removed for three different boxes where measured the standard deviation of the map. The effect of PCA on high orbital velocity residuals is less than that on low velocity residuals, so calculating the standard deviation in the red box will underestimate the S/N. Conversely, the effect of PCA on low velocity residuals is much stronger, so the residuals will be smaller than those for high velocity, and the S/N will be overestimated. We then decided to calculate the standard deviation from the blue box, which mitigates this effect by taking the residuals at each velocity, but still far from the RM residuals or planetary signals, to avoid misinterpreting the standard deviation of the residual. The same method was used for each species observed.
Several maps, such as Fe, Mg, Ca, or Cr, show a parasitic signal at km/s and km/s. This signal is a residual of the Rossiter-McLaughlin effect observed in the CCF map in Fig. 5. The Rossiter-McLaughlin effect presents an anti-correlation signal at negative . Therefore, it is not shown here, but the residual positive correlation visible in yellow around the anti-correlation signal in the CCF map is the parasite signal observed in the diagram mentioned previously. Possible biases due to Rossiter Mc-Laughing effect are present at phase below -5 degree and are represented in grey on the trail map of Fig. 7. The positive positive correlation at phases above 5 degree is far from the planetary signal 5 ( km/s). For the case of Mg, a signal is also visible at expected but = 100 km.s-1. This is due to a strong Fe+ line in the Mg triplet. This is also visible in the CCF maps of both species, where the residual signal of the other can be observed as highlighted in Fig. 18.
Fig. 20 highlights some limitations of the Gaussian profiles used to parameterize the iron trail of TOI-1518 b as most one-dimensional CCF do not follow a simple Gaussian profile. Then we decided to center the fitted Gaussian profile on the maximum of the 1D-CCF even if this resulted in a misestimation of the FWHM and probably of the error bars of the measured V. Another uncertainty might be due to the PCA applied to the data. Fig. 19 presents the iron trail of TOI-1518 b with different numbers of principal components removed. The square root of the variance of the Doppler shifts across all numbers of removed components, PCA has been added to the uncertainty quoted in the covariance matrix of the Gaussian fit obtained from scipy.optimize.curve fit. The PCA does not change the transit trend even when many components are removed for all species except Ca+ (Fig.21). The PCA is essential for detecting faint species. To maintain consistency, we removed the same number of components (three) from all species, even for those where it might not have been necessary. The only exception is Ca+, as the PCA significantly impacts the signal, even when fewer components are removed, due to its very high signal strength. Therefore, we decided not to apply PCA to Ca+ in order to achieve a more robust analysis of its trail. To understand the physics behind these six time-resolved absorption, a comparison with Global Circulation Models (GCMs) is conducted in Sect. 4.
4 Comparison with Global Circulation Models
The 3D nature of UHJs can explain the trace of iron in the CCF map observed in this work. We compare the iron signal’s blueshift with two different 3D global circulation models (GCMs) scenarios to interpret the data.
4.1 Model description
We consider five SPARC/MITgcm models of TOI 1518 b, three are presented in Fig. 9. The SPARC/MITgcm was initially introduced by Showman et al. (2009).
It has been widely used to study the atmospheric physics and chemistry of (ultra-)hot Jupiters (Fortney et al., 2010; Showman et al., 2013; Kataria et al., 2013; Parmentier et al., 2018; Tan et al., 2024). Our models of TOI-1518 b are based on work by Tan et al. (2024). We explore the effect of atmospheric drag (Showman et al., 2013; Komacek & Showman, 2016; Parmentier & Crossfield, 2018), in considering four models with drag timescales from s (strong drag) s (weak drag). The drag timescale encompasses a wide range of physical mechanisms, including turbulent mixing (Li & Goodman, 2010), Lorentz-force braking of ionized winds within the planet’s magnetic field (Perna et al., 2010a), and Ohmic dissipation (Perna et al., 2010b).
These GCMs account for heat transport due to H2 dissociation and recombination (e.g., Bell & Cowan (2018); Komacek & Tan (2018);Tan & Komacek (2019); Roth et al. (2021)). H2 thermally dissociates on the dayside, after which atomic hydrogen gets advected to the nightside, where it recombines into H2 and releases latent heat. When the atmospheric circulation is predominantly eastward, most of this heat is dumped on the evening limb, resulting in a temperature asymmetry between the eastern and western regions of the atmosphere (first and second column in Fig.9).
Drag restores energy to the atmosphere. Increasing drag strength (i.e lowering ) slows down winds in the atmosphere hindering this model’s heat transport. In strong drag cases, the temperature structure becomes symmetric, resulting in similar chemical compositions in the morning and evening limbs. (last column in Fig. 9). In the strong-drag model, there is only a day-to-night flow as the equatorial jet gets suppressed. Table 4 summarizes some other important parameters of the two SPARC/MITgcm models. We refer to Table 1 in Tan et al. (2024) for the full list of opacities considered in their radiative transfer. All models were run at a horizontal resolution of C32, corresponding to roughly 128 cells in longitude and 64 in latitude. Before computing phase-dependent spectra of the GCMs with gCMCRT, we bin the outputs down to 32 latitudes and 64 longitudes, as in Wardenier et al. (2021, 2023, 2024).
| Parameter | Value |
| Orbital Period | 1,6442 s (1.903 days) |
| Pressure range | 200 2 |
| Radius at bottom | 1.3405 108m (1,875 R) |
| Gravity | 10.56 m/s2 |
| Horizontal resolution | C32 |
| Vertical resolution | 53 layers |
| Metallicity and C/O | 1 solar |
| H/H2 heat transport | {,,,,} |
| Drag timescale | {,s,s,s,s} |
| Radiative transfer | non-grey (see ) |
4.2 Injection and Cross-Correlation maps
We compute phase-dependent transmission spectra of the two GCM models across the MAROON-X spectral range (between 490 and 920 nm) using gCMCRT (Lee et al., 2022). The calculations account for Doppler shifts due to planet rotation and winds. Section 2 of Wardenier et al. (2023) shows the radiative transfer and post-processing details. Therefore, we will only briefly summarize the gCMCRT setup for TOI-1518 b analysis. Before feeding the GCM outputs into gCMCRT, we map the atmospheric structures onto a 3D grid with altitude (instead of pressure) as a vertical coordinate. We account for the fact that each atmospheric column has a different scale height set by local gravity, temperature, and mean molecular weight. For each of the five TOI-1518 b models, we simulate 25 spectra (equidistant in orbital phase) between phase angles 8 degrees, covering the in-transit part of our observations. We assume an orbital impact parameter of 0.9036, a semi-major axis of 0.039 au, a stellar radius of 1.95 R⊙, and an orbital period of 1.90 days. At each orbital phase angle, gCMCRT simulates a transmission spectrum by randomly shooting photon packets at the part of the planet limb that is blocking the star and evaluating the optical depth encountered by each photon packet. Due to the geometry of TOI-1518 b (Fig. 16), the illumination differs from the case of WASP-121 b where the impact parameter is near 0 (Fig. 10, Right Panel). In this calculation, the code accounts for Doppler shifts imparted on the opacities by the radial component of the local wind vector and planet rotation (Wardenier et al., 2021). The transit depth at a specific wavelength is calculated by averaging it over all photon packets. To accurately represent the shapes, depths, and shifts of the spectral lines, we use 105 photon packets per wavelength. Since we do not explicitly account for scattering, the direction of propagation for the photon packets remains constant throughout the calculation. The spectra are calculated at a resolution of , which differs from Wardenier et al. (2023, 2024). We include the same set of continuum opacities and line species for the radiative transfer as in Wardenier et al. (2023). To properly compare the GCM spectra with the data, we need to inject the GCMs into the data and perform PCA, similar to what we did for the data. The different steps were as follows:
Inject the spectra at each orbital phase of the observations at the expected Doppler shift calculated with the expected of Equation 1.
Perform PCA on the data+GCM removing three components, the same number of components as for the data in the case of iron.
Subtract the post-PCA data from the combined data+GCM to isolate the GCM signal post-PCA and remove the noise.
Do the cross-correlation with the same Fe template and methods used for Sect. 3.1.
The resulting CCFs for the case of the drag model are shown in Fig. 10 . The difference between the two models at different impact parameters b is also represented with the case of b = 0 on the upper row and b=0.9 on the lower row.
Figs. 11 and 12 show the CCF signals of Fe and Fe+ that we obtain for each of the models, with the real data plotted on top. We interpolated the cross-correlation results over the same phase grid as for the observations of the two transits. We then performed the same analysis using the fit with a Gaussian profile, which allowed us to extract the position, amplitude, and FWHM of the model’s signal.
4.3 Discussion
Fig. 10 shows the importance of considering the impact parameter in calculating the spectra. We clearly see a big difference in the spectrum for the same GCM output of the SPARC/MITgcm. The blueshift is way more pronounced in the case of a high-impact parameter. The jump in the signal driven by planet rotation also occurs earlier in the case of a high-impact parameter (around phase = -2.5 deg) than in the case of impact parameter b =0, where it happens at mid-transit. This is because the illuminated part of the planet differs in the two cases. When b=0, the leading limb is illuminated at the start of the transit, both limbs are illuminated in the middle, and the trailing limb is illuminated at the end. In contrast, only the southern parts of the limbs are illuminated in the case of b =0.9.
Fig.11 (Left Panel) shows the Fe trails of the five GCMs simulations, along with the Doppler shifts of Fe measured in section 3.1. The absorption trails stand between the stronger drag models ( and s). The weak drag and drag-free models produce Doppler shifts that are too much blueshifted due to strong winds that are not attenuated by drag effects.
Fig.12 now compares the GCM outputs to the Fe+ trace. There is a pretty good match between the observation and the stronger drag model (s in purple), especially since the GCMs are not optimized to fit the data. Overall, the fact that both Fe and Fe+ trails are very similar points towards a strong drag scenario, where the iron is fully condensed on the nightside and both Fe and Fe+ probe similar regions close to the limb.
Fig.13 shows the trails for the 6 species of Fig.7 with the signals of the stronger drag models. For Ca, Na, Mg, and Ca+ the signal straddles between both these GCMs, as in the case of Fe. The strong drag scenario is then preferred also by the other species detected
However the distribution of ionized iron (Fe+), which is present on the entire day side, and neutral iron (Fe), found only at the limbs and in the inner part (Fig.9), may account for the variations in the models and data shape of the trail. In particular, the drag-free and weak drag models exhibit a Fe signal influenced primarily by strong wind effects, with velocities ranging from around -5 to -12 km/s. The wind in the two stronger drag models is almost nullified at the Fe localization. There is a clear difference between the ingress and egress parts of the transit, which is caused by the redshift and blueshift of the planet’s rotation (Fig.11). The wind at the location of Fe+ in the upper atmosphere is less affected in the strong drag model, leading to a more continuous trail between ingress and egress part of the transit (Fig.12).
This effect is more extreme in the case of the Ca+ signal, which is redshifted over compared to the models and the other trails. Ionised species appear then more redshifted than neutral species. The models used in this study assume consistent drag in the atmosphere; however, this finding highlights the potential variation of drag with altitude. We expect neutral species to dominate in the lower part of the atmosphere, while ionized species will be more prevalent in the upper layers. This variation in drag could be attributed to an increase in magnetic drag with altitude, as demonstrated in previous studies (e.g., Beltz et al. 2022).
The agreement with an intense drag model atmosphere was already observed in the case of WASP-121 b by Wardenier et al. 2024 by comparing with ESPRESSO data at with similar models. Moreover, as in WASP-76 b (e.g., Wardenier et al. 2021; Savel et al. 2022), the Fe trail of TOI-1518 b provides clear evidence of a strong thermochemical asymmetry between the morning and evening terminator. Fig. 11, we can see that in the first half of the transit, the data are more blueshifted than the models with stronger drags (between phases at -5 and 0 degrees), which probably implies that the signal from the trailing limb is stronger (compared to a model with no asymmetry)
Then, a similar scenario to the one observed in WASP-76 b (Fig. 2 in Wardenier et al. 2021) seems to appear where, at the beginning of the transit, the Fe signal is mainly influenced by the leading limb, where the redshift from the planet’s rotation balances out the blueshift from day-to-night winds. The Doppler shift should become more negative as the trailing limb comes into view because the planet’s rotation and day-to-night winds cause a blueshift in the signal.
Looking at Fig. 11, the Doppler-shift measurement of iron (Fe) in the third bin of the transit (just before mid-transit) may be considered an outlier, but this could also be attributed to a limitation of the Gaussian profile used in this context. In Fig. 10, the CCF of the GCM exhibits a double-peak structure with one peak near 0 km/s and another near -7 km/s. The data signal does not seem to reflect this double-peak scenario, instead showing a non-Gaussian signal, as discussed in Sect. 3.4 (see Fig. 20).
The amplitude of the signal shown in the middle panel of Fig.11 and Fig.12 is mainly influenced by the portion of the planet that transits in front of the star, depicted by the yellow dashed line. All models follow this trend even though it is more robust in the second part of the transit compared to the data. The signal amplitude does not underline any asymmetry between the two limbs or between the models because of the specific geometry of TOI-1518 b. It still interesting to note that the amplitude of the signal in the Fe simulations is lowering with stronger drag effect due to the double peak scenarios underlined before. This double peak scenario appear because of wind reduced by the drag as discussed above. For the Fe+ simulations, the winds is higher and so the signal is smoother with all the models following the same trend due to the fraction of the planet.
The FWHM is also in agreement with all models in the first order. This parameter is not constrained by the approach of a simple Gaussian fit as presented in Fig.20, and another approach would be more efficient, but this is not the primary purpose of this paper. The fact that the models have FWHM less important with drag effect increasing may be due to the double-peak structure present from phase -5 degrees to +2.5 degrees, corresponding to the lower FWHM observed. As our simple Gaussian profile only fits one of the two peaks, there may be an underestimation of the FWHM of the model signal compared to the data signal, which does not emphasise this double peak shape.
5 Retrieval analysis
After detecting the species shown in Sect. 3.1 thanks to cross-correlation analysis, the next step is to explore the abundances of these species in comparison to solar values (This comparison is possible due to the solar metallicity of the host star). One approach developed in Brogi & Line 2019 is to use a Bayesian atmospheric retrieval framework with high-resolution cross-correlation spectroscopy (HRCCS) that relies on the cross-correlation between data and models for extracting the planetary spectral signal. This approach permits the characterization of many atmospheres of UHJs and puts constraints on abundances (Line et al., 2021; Kasper et al., 2021, 2023; Brogi et al., 2023).
5.1 CHIMERA codes
Following the above method, we applied the Brogi & Line 2019 cross-correlation-to-log-likelihood retrieval framework to derive the molecular volume mixing ratios and the temperature layer we are probing. For the retrieval process, we used the basic CHIMERA “free-retrieval” (Line et al., 2013; Kreidberg et al., 2015) paradigm, which assumes constant-with-altitude gas mixing ratios and uses a simple isothermal T-P profile. Pelletier et al. 2023 shows that a simple isothermal profile would give a similar estimation of chemical abundance ratio to a more complex T-P profile. The retrieval parameters specific to our analysis and their prior ranges are provided in Table 5. A more detailed description of the high-resolution GPU-based radiative transfer method and loglikelihood implementation within pymultinest (Feroz et al., 2009; Buchner et al., 2014) is given in Line et al. 2021. Retrieval parameters and their priors are reported in Table 5.
| Parameter | Description | Prior |
| T0 | Isothermal temperature | 1500 - 4500 K |
| xRp | Scaled radius of the planet | 0.5 - 1.5 |
| Mp | Mass of the planet | 0.5 - 2.3 |
| Log (Pc) | Continuum | -6 - 0 |
| Kp | Planet Orbital velocity | 100 - 300 km/s |
| Vsys | Systemic Velocity | -100 - 100 km/s |
| Log(a) | Model scaling factor | -2 - 2 |
| Log(H-) | Continuum | -12 - 0 |
| Log(Fe) | log gas volume mixing ratio | -12 - 0 |
| Log(/Fe) | Relative abundance | -8 - 1 |
| Log() | Ionised fraction (for Fe,Ti,Ca) | -2 - 0 |
5.2 Results of the retrieval analysis
We combined the blue and red arm MAROON-X data for our retrieval for both transits. We included most of the species detected in Sect. 3.1 TiO and abundance proxies for the H bound-free and free-free continua. For detected species in both ionized and neutral states, we compute an ionized fraction of them. (Example with Fe in equation 2):
Because we only have a 3-sigma upper limit of 2.3 M for the planetary mass (Cabot et al., 2021), we include the mass as a free parameter that spanned a range between maximum Mp = 2.3 M and Mp = 0.5 M. For this analysis, we consider the two detectors independently as for the CCF, and we add the likelihood from the red and the blue detector analysis and then from both transit. For the retrieval analysis, we used three number of principal components as for the CCF analysis. Because we had to put the same number of principal component for each species, the Ca+ lines may be affected a bit by the PCA. The results are summarized in Fig. 14 and Fig. 15 and the whole corner plot is given in the Appendix (Fig. 24).
5.3 Discussion of the retrieval analysis
The and Vres distribution of the retrieval analysis (Fig. 14) are consistent with the iron signal detected in section 3.1. The measured Vres () and () also agree with the detection of Cabot et al. (2021) ( km/s and = km/s). We observed that the velocity parameters obtained are influenced predominantly by the strongest absorber, particularly Fe. Other species may have minor differences in and , as discussed in section 3.1. However, we have chosen not to individually adjust and for each species to derive meaningful abundance ratios. Some species detected in the Cross-Correlation analysis were not included in the retrieval framework due to a lack of opacities files like Ba+, Si, Mn, and the HyVO line list for VO (for which we used the VOmyt line list instead).
Our retrievals converge towards values of Mg, Ca, and Fe abundances on TOI-1518 b that are consistent with proto-solar abundances. Cr, Ti, and V, however, differ at the 1 level. The lower abundances of Cr, V and Ti can plausibly be explained by their condensation on the planet nightside or by the fact that they are bound to other molecules (e.g. TiO, VO), or by the fact that they are partially ionized ( e.g. Na+, Cr+). We already showed that the line lists are essential for the cross-correlation (Fig. 8). We could not use the HyVO line list in the retrieval analysis. The VOmyt line list leads to a nondetection of VO in the retrieval analysis, as in the CCF analysis. The missing vanadium is then probably present in VO or is ionized. For titanium, the subsolar value could also be explained by the non-detection of TiO. Further observations with bluer or redder instruments to detect molecules or ionized species are needed to determine why such elements are subsolar.
Even if the isothermal TP profile could bias our results, Pelletier et al. 2023 showed that different temperature-pressure profiles inferred similar elemental abundance ratios.
6 Conclusions
This study presents an in-depth analysis of TOI-1518 b, an ultra-hot Jupiter. It uses transit observations, from MAROON-X on Gemini-N, to explore its atmospheric dynamics and chemical composition. Our findings offer new insights into the unique characteristics of this extreme exoplanet.
Our cross-correlation analysis focused on detecting atomic and molecular species within the atmosphere. We report the detection of 14 different species. High-resolution spectroscopy allowed us to identify ionized and neutral species with ionized metals such as Fe+, Ca+, and Ti+. The detection of these ionized species, as opposed to their neutral counterparts, underscores the extreme temperatures of TOI-1518 b’s atmosphere, where thermal ionization is significant. Additionally, we reported the detection of vanadium oxide (VO), a critical absorber that has implications for thermal inversions in ultra-hot Jupiters. The presence of VO adds a crucial piece to the puzzle of understanding the thermal structure and chemical processes in such extreme environments. This was possible thanks to the newer HyVO line list that sould be used in other studies were detection of VO failed with previous line lists
We investigated the atmospheric wind dynamics by analyzing the blueshift of multiple species in the observed spectra as a function of the planet’s orbital phase. The blueshift of iron is consistent with previous observations of other ultra-hot Jupiters. By combining the signal from different species, particularly Fe+ and Fe, and comparing them with Global Circulation Models, we conclude that a strong drag is needed (. Furthermore, ionized species need more substantial drags compared to neutral species, probably because of the amplified magnetic effects present in the upper atmosphere.
The retrieval analysis provided constraints on the abundance of various chemical species in the atmosphere. Our results suggest non-solar elemental abundance ratios for Vanadium, Magnesium, Chromium, and Titanium relative to iron, highlighting the complex chemistry of TOI-1518 b.
Overall, this study demonstrates the power of combining high-resolution spectroscopy with advanced modeling techniques to probe the atmospheres of ultra-hot Jupiters. They shed light on the specific properties of TOI-1518 b and contribute to the broader understanding of atmospheric dynamics and chemistry in these extreme exoplanets. Future studies should continue to refine these models and expand observational efforts to explore the diversity and complexity of ultra-hot Jupiter atmospheres.
Acknowledgements.
This project has received funding from Agence Nationale de la Recherche (ANR) for the grant ANR EXOWINDS.This work was supported by the French government through the France 2030 investment plan managed by the National Research Agency (ANR), as part of the Initiative of Excellence Université Côte d’Azur under reference number ANR-15-IDEX-01. The authors are grateful to the Université Côte d’Azur’s Center for High-Performance Computing (OPAL infrastructure) for providing resources and support.
J.P.W. acknowledges support from the Trottier Family Foundation via the Trottier Postdoctoral Fellowship.
M.R.L. and J.L.B. acknowledge support from NASA XRP grant 80NSSC19K0293 and NSF grant AST-2307177. J.L.B. acknowledges funding for the MAROON-X project from the David and Lucile Packard Foundation, the Heising-Simons Foundation, the Gordon and Betty Moore Foundation, the Gemini Observatory, the NSF (award number 2108465), and NASA (grant number 80NSSC22K0117).
Appendix A Cross-Correlation-Function
Appendix B Global Circulation Models
Appendix C Retrieval analysis